OpenStudy (anonymous):

x^2+y^2=14 x+y=7 What is x-y ?

OpenStudy (anonymous):

(7-y)^2+y^2=14

OpenStudy (anonymous):

y^2-14y+49+y^2=14

OpenStudy (anonymous):

\[(x+y)^2 = 7^2 \iff x^2+y^2+2xy = 49 \iff 14+2xy = 49 \] \[2xy = 35 \iff xy = 17.5\] \[(x-y)^2 = x^2+y^2-2xy =14-2(17.5) = -21\] trying to be snazzy and something doesnt look right lolol

OpenStudy (anonymous):

2y^2-28y+49=0

OpenStudy (anonymous):

joemath i need ur help

OpenStudy (anonymous):

canu try helping her

OpenStudy (anonymous):

agree

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

i was following along until you had a perfect square=-21

OpenStudy (anonymous):

thts easy hold up

OpenStudy (anonymous):

lets imagine it is true that \[x-y=\sqrt{21}i\] and if this makes sense.

OpenStudy (anonymous):

now i am wondering whether there was a typo and it should be \[x^2-y^2=14, x+y=7\] then you get the answer in a nanosecond